OpenStudy (anonymous):

Find the Domain of f(x)= e^√((x+2)/(x^2-1)) My working is as follows f(x)= e^√((x+2)/(x^2-1)) √((x+2)/(x^2-1))≥0 x^2-1>0 x^2>1 x>1 D:All real numbers>1 However when I graph the function, x can be equal to -2?

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

your job it to solve \[\frac{x+2}{x^2-1}\geq 0\]

OpenStudy (anonymous):

two mistakes here in your calculation you ignored the \(x+2\)in the numerator, and \[x^2-1>0\\x^2>1\\x>1\] is not correct

OpenStudy (anonymous):

the second mistake is less important than the first one, since you have to solve \[\frac{x+2}{x^2-1}\geq 0\] in any case

OpenStudy (anonymous):

I think I understand what you're saying, thank you.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Could you elaborate a little further on the second mistake, how would I find the domain of x^2-1>0?

OpenStudy (anonymous):

Also if my \[x \ge 2\] then my denomiator could be equal to zero?

OpenStudy (anonymous):

-2*